# Localized excitation on the Jacobi elliptic periodic background for the (n+1)-dimensional generalized Kadomtsev-Petviashvili equation

### v1

@@ 5.1 First order breather waves on the traveling Jacobi elliptic function periodic wave background
for $i=2,3,\cdots,n$. $\beta,\delta,\omega_{i},\sigma_{i}\in\mathbb{R}$ are arbitrary, and $\phi_{\gamma}\in[0,\frac{\pi}{2}]$. The appropriate normalized phase shift for the background Jacobi elliptic wave is expressed as follows:
@@
$F_{2}=\left(Z_{x_{1}}(\eta-\alpha)-Z_{x_{1}}(\eta+\alpha)\right)\Theta^{2}(\eta-\alpha)$
×$\qquad\ \times\mathrm{exp}(2\kappa_{1})+\left((Z(\eta+\alpha)-Z(\eta-\alpha)-2\mu_{1})^{2}\right.$
$\qquad\ \left.+Z_{x_{1}}(\eta-\alpha)-Z_{x_{1}}(\eta+\alpha)\right)\Theta(\eta+\alpha)\Theta(\eta-\alpha),$
@@
Figure 2 illustrates the evolution and corresponding contour plots for the solution (62) in the space of different spatial variables $x_{1}$, $x_{2}$, and $x_{3}$ and time variable $t$ under some specified parameter values. Figure 2(a) demonstrates the structure of dark breather waves on the periodic background, moving at a speed of $c$ in the $x_{1}$-$t$ axis, alongside the dark breather exhibiting an inverse width $\mu_{2}$ and a speed of $c_{2}$. This breather is created with a phase shift of $-2\alpha$. Similarly, in both Figure 2(b) and Figure 2(c), the periodic background wave moves in the direction of $t$. The inverse width of the dark breather, denoted as $\omega_{i}(i=2,~{}3)$, exhibits a phase shift of $-2\alpha$, and its propagation velocity is $c_{2}/p_{i}~{}(i=2,~{}3)$, in that order. Obviously, other parameters can be chosen to represent the evolution and dynamics of solution (62) in $x_{i}$-$t$ space, and provide theoretical explanations for its corresponding physical reality.
In summary, spectrum analysis reveals that on the Jacobi elliptic function periodic wave background (19), when $\lambda$ falls within $[\lambda_{1+},+\infty)$, the gKP equation (1) can exhibit bright breather waves, whereas for $\lambda$ within $[\lambda_{3+},\lambda_{2+}]$, the equation shows dark breather waves.
@@ 5.2 Second order breather waves on the traveling Jacobi elliptic function periodic wave background
where $\bar{\varphi}=\mathrm{Wr}[\varphi_{1},\varphi_{2}]$. Based on the conclusion of Corollary 3, in order to facilitate the analysis of the dynamic behavior of the breather wave (63), the cases of $\lambda$ in the two intervals of $[\lambda_{1+},+\infty)$ and $[\lambda_{3+},\lambda_{2+}]$ are discussed, respectively.
@@
×$\quad\ \ \times\Theta(\eta+\gamma_{1})\Theta(\eta+\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta+\gamma_{1})\Theta(\eta-\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta-\gamma_{1})\Theta(\eta+\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta-\gamma_{1})\Theta(\eta-\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}),$
@@
$\bar{\mu}_{j}=\frac{\sqrt{\lambda_{j}+1+\delta\omega_{2}+k^{2}}\sqrt{\lambda_{j}+\delta\omega_{2}+k^{2}}}{\sqrt{\lambda_{j}+1+\delta\omega_{2}}}$
$\quad\ \ \ -Z(\phi_{\gamma},k),$
@@
×$\quad\ \ \ \times\sqrt{\lambda_{j}+1+\delta\omega_{2}}-\sum_{i=2}^{n}\sigma_{i}\omega_{i}\big{)}/\bar{\mu}_{j},$
@@
$\bar{F}=\left(Z_{x_{1}}(\eta+\gamma_{2})\left(3Z(\eta+\gamma_{2})+3\bar{\mu}_{2}+Z(\eta+\gamma_{1})\right.\right.$
$\quad\ \ \left.\left.+\bar{\mu}_{1}\right)-Z_{x_{1}}(\eta+\gamma_{1})\left(3Z(\eta+\gamma_{1})+3\bar{\mu}_{1}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta+\gamma_{2})+\bar{\mu}_{2}\right)+\left(Z(\eta+\gamma_{2})-Z(\eta+\gamma_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta+\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta-\gamma_{2})\left(3Z(\eta-\gamma_{2})+3\bar{\mu}_{1}+Z(\eta+\gamma_{1})\right.\right.$
$\quad\ \ \left.\left.-\bar{\mu}_{2}\right)-Z_{x_{1}}(\eta+\gamma_{1})\left(3Z(\eta+\gamma_{1})-3\bar{\mu}_{2}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta-\gamma_{2})+\bar{\mu}_{1}\right)+\left(Z(\eta-\gamma_{2})-Z(\eta+\gamma_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta-\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta+\gamma_{2})\left(3Z(\eta+\gamma_{2})-3\bar{\mu}_{1}+Z(\eta-\gamma_{1})\right.\right.$
$\quad\ \ \left.\left.+\bar{\mu}_{2}\right)-Z_{x_{1}}(\eta-\gamma_{1})\left(3Z(\eta-\gamma_{1})+3\bar{\mu}_{2}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta+\gamma_{2})-\bar{\mu}_{1}\right)+\left(Z(\eta+\gamma_{2})-Z(\eta-\gamma_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta+\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta-\gamma_{2})\left(3Z(\eta-\gamma_{2})-3\bar{\mu}_{2}+Z(\eta+\gamma_{1})\right.\right.$
$\quad\ \ \left.\left.-\bar{\mu}_{1}\right)-Z_{x_{1}}(\eta-\gamma_{1})\left(3Z(\eta-\gamma_{1})-3\bar{\mu}_{1}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta-\gamma_{2})-\bar{\mu}_{2}\right)+\left(Z(\eta-\gamma_{2})-Z(\eta-\gamma_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta-\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}),$
$\bar{G}=\left(\left(Z(\eta+\gamma_{2})-Z(\eta+\gamma_{1})-\bar{\mu}_{1}+\bar{\mu}_{2}\right)\left(Z(\eta+\gamma_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta-\gamma_{2})-Z(\eta+\gamma_{1})+\bar{\mu}_{1}+\bar{\mu}_{2}\right)\left(Z(\eta+\gamma_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta+\gamma_{2})-Z(\eta-\gamma_{1})+\bar{\mu}_{1}+\bar{\mu}_{2}\right)\left(Z(\eta-\gamma_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta-\gamma_{2})-Z(\eta-\gamma_{1})-\bar{\mu}_{1}+\bar{\mu}_{2}\right)\left(Z(\eta-\gamma_{1})\right.\right.$
@@
[…]
(a)                                              (b)                                              (c)
Figure 3: Propagation and contour plots for nonlinear wave solution (65) via two-times DT, where $k=0.3$, $\lambda_{1}=1.8$, $\lambda_{2}=1.4$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, $\omega_{i}(i=4,\cdots,n)=0$. (a) $x_{2}=x_{3}=1$; (b) $x_{1}=x_{3}=1$; (c) $x_{1}=x_{2}=1$.
(a)                                              (b)                                              (c)
Figure 4: Propagation and contour plots for nonlinear wave solution (67) via two-times DT, where $k=0.6$, $\lambda_{1}=-2.1$, $\lambda_{2}=-2.3$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, $\omega_{i}(i=4,\cdots,n)=0$. (a) $x_{2}=x_{3}=1$; (b) $x_{1}=x_{3}=1$; (c) $x_{1}=x_{2}=1$.
@@
×$\quad\ \ \times\Theta(\eta+\alpha_{1})\Theta(\eta+\alpha_{2})\mathrm{exp}(-\hat{\kappa}_{1}-\hat{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta+\alpha_{1})\Theta(\eta-\alpha_{2})\mathrm{exp}(-\hat{\kappa}_{1}+\hat{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta-\alpha_{1})\Theta(\eta+\alpha_{2})\mathrm{exp}(\hat{\kappa}_{1}-\hat{\kappa}_{2})$
@@
×$\quad\ \ \times\Theta(\eta-\alpha_{1})\Theta(\eta-\alpha_{2})\mathrm{exp}(\hat{\kappa}_{1}+\hat{\kappa}_{2}),$
@@
×$\quad\ \ \ \left.\times\sqrt{\lambda_{j}+1+\delta\omega_{2}}-\sum_{i=2}^{n}\sigma_{i}\omega_{i}\right)/\hat{\mu}_{j},$
@@
$\hat{F}=\left(Z_{x_{1}}(\eta+\alpha_{2})\left(3Z(\eta+\alpha_{2})-3\bar{\mu}_{2}+Z(\eta+\alpha_{1})\right.\right.$
$\quad\ \ \left.\left.+\bar{\mu}_{1}\right)-Z_{x_{1}}(\eta+\alpha_{1})\left(3Z(\eta+\alpha_{1})-3\bar{\mu}_{1}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta+\alpha_{2})-\bar{\mu}_{2}\right)+\left(Z(\eta+\alpha_{2})-Z(\eta+\alpha_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta+\alpha_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta-\alpha_{2})\left(3Z(\eta-\alpha_{2})-3\bar{\mu}_{1}+Z(\eta+\alpha_{1})\right.\right.$
$\quad\ \ \left.\left.+\bar{\mu}_{2}\right)-Z_{x_{1}}(\eta+\alpha_{1})\left(3Z(\eta+\alpha_{1})+3\bar{\mu}_{2}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta-\alpha_{2})-\bar{\mu}_{1}\right)+\left(Z(\eta-\alpha_{2})-Z(\eta+\alpha_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta-\alpha_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta+\alpha_{2})\left(3Z(\eta+\alpha_{2})+3\bar{\mu}_{1}+Z(\eta-\alpha_{1})\right.\right.$
$\quad\ \ \left.\left.-\bar{\mu}_{2}\right)-Z_{x_{1}}(\eta-\alpha_{1})\left(3Z(\eta-\alpha_{1})-3\bar{\mu}_{2}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta+\alpha_{2})+\bar{\mu}_{1}\right)+\left(Z(\eta+\alpha_{2})-Z(\eta-\alpha_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta+\alpha_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
$\quad\ \ +\left(Z_{x_{1}}(\eta-\alpha_{2})\left(3Z(\eta-\alpha_{2})+3\bar{\mu}_{2}+Z(\eta+\alpha_{1})\right.\right.$
$\quad\ \ \left.\left.+\bar{\mu}_{1}\right)-Z_{x_{1}}(\eta-\alpha_{1})\left(3Z(\eta-\alpha_{1})+3\bar{\mu}_{1}\right.\right.$
$\quad\ \ \left.\left.+Z(\eta-\alpha_{2})+\bar{\mu}_{2}\right)+\left(Z(\eta-\alpha_{2})-Z(\eta-\alpha_{1})\right.\right.$
@@
×$\quad\ \ \times\Theta(\eta-\alpha_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2}),$
$\hat{G}=\left(\left(Z(\eta+\alpha_{2})-Z(\eta+\alpha_{1})-\hat{\mu}_{1}+\hat{\mu}_{2}\right)\left(Z(\eta+\alpha_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta-\alpha_{2})-Z(\eta+\alpha_{1})+\hat{\mu}_{1}+\hat{\mu}_{2}\right)\left(Z(\eta+\alpha_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta+\alpha_{2})-Z(\eta-\alpha_{1})+\hat{\mu}_{1}+\hat{\mu}_{2}\right)\left(Z(\eta-\alpha_{1})\right.\right.$
@@
$\quad\ \ +\left(\left(Z(\eta-\alpha_{2})-Z(\eta-\alpha_{1})-\hat{\mu}_{1}+\hat{\mu}_{2}\right)\left(Z(\eta-\alpha_{1})\right.\right.$
@@
[…]

### v2

@@ 5.1 First order breather waves on the traveling Jacobi elliptic function periodic wave background
for $i=2,3,\cdots,n$. $\beta,\delta,\omega_{i},\sigma_{i}\in\mathbb{R}$ are arbitrary, $D$ is the dispersion parameter and $\phi_{\gamma}\in[0,\frac{\pi}{2}]$. The appropriate normalized phase shift for the background Jacobi elliptic wave is expressed as follows:
@@
$F_{2}=\left(Z_{x_{1}}(\eta-\alpha)-Z_{x_{1}}(\eta+\alpha)\right)\Theta^{2}(\eta-\alpha)\mathrm{exp}(2\kappa_{1})$
$\qquad\ +\left((Z(\eta+\alpha)-Z(\eta-\alpha)-2\mu_{1})^{2}+Z_{x_{1}}(\eta-\alpha)\right$.
$\qquad\ \left.-Z_{x_{1}}(\eta+\alpha)\right)\Theta(\eta+\alpha)\Theta(\eta-\alpha),$
@@
It is observed from the explicit form of solution (62) that the linear dispersion coefficients enter the dark breather dynamics through the propagation velocity. In particular, the velocity parameter $c_{2}$ appearing in (60) depends on the effective dispersion parameter $D=-\sum_{i=2}^{n}\sigma_{i}\omega_{i}$, which indicates that linear dispersion also affects the longitudinal propagation of dark breather waves. This dependence provides a direct interpretation of how dispersion parameters are reflected in the dynamical behavior of the solution, while the overall breather structure remains unchanged.
Figure 2 presents the evolution profiles and corresponding contour plots of the solution (62) in the $x_{1}$$t$, $x_{2}$$t$, and $x_{3}$$t$ planes under a specified parameter configuration. As shown in Fig. 2(a), a dark breather structure is formed on the same periodic background, propagating with velocity $c$ along the $x_{1}$$t$ direction. The dark breather is characterized by an inverse width $\mu_{2}$, a propagation velocity $c_{2}$, and a phase shift of $2\alpha$. Likewise, Figs. 2(b) and 2(c) demonstrate the breather dynamics in the $x_{2}$$t$ and $x_{3}$$t$ planes, respectively. The inverse widths are given by $\omega_{i}~(i=2,3)$, while the corresponding propagation velocities take the form $c_{2}/p_{i}~(i=2,3)$, both sharing the same phase shift $2\alpha$. Different choices of parameters in solution (62) allow one to describe a variety of dark breather evolution scenarios, which are relevant to different physical realizations.
Finally, spectral analysis indicates that on the Jacobi elliptic function periodic background (19), the nature of breather solutions is determined by the location of the spectral parameter $\lambda$. Specifically, when $\lambda\in[\lambda_{1+},+\infty)$, the gKP equation (1) admits bright breather waves, whereas for $\lambda\in[\lambda_{3+},\lambda_{2+}]$, dark breather waves are generated.
@@ 5.2 Second order breather waves on the traveling Jacobi elliptic function periodic wave background
where $\bar{\varphi}=\mathrm{Wr}[\varphi_{1},\varphi_{2}]$. Based on the conclusion of Corollary 3, in order to facilitate the analysis of the dynamic behavior of the breather wave (63), the cases of $\lambda$ in the two intervals of $\lambda_{1+},+\infty)$ and $\lambda_{3+},\lambda_{2+}]$ are discussed, respectively.
@@
$\quad\ \ \times\Theta(\eta+\gamma_{1})\Theta(\eta+\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta+\gamma_{1})\Theta(\eta-\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta-\gamma_{1})\Theta(\eta+\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta-\gamma_{1})\Theta(\eta-\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}),$
@@
$\bar{\mu}_{j}=\frac{\sqrt{\lambda_{j}+1+\delta\omega_{2}+k^{2}}\sqrt{\lambda_{j}+\delta\omega_{2}+k^{2}}}{\sqrt{\lambda_{j}+1+\delta\omega_{2}}}-Z(\phi_{\gamma},k),$
@@
$\quad\ \ \ \times\sqrt{\lambda_{j}+1+\delta\omega_{2}}+D\big)/\bar{\mu}_{j},$
@@
[…]
@@
$\quad\ \ \times\Theta(\eta+\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta-\gamma_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta+\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta-\gamma_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2}),$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
The 3D and corresponding contour plots of second order breather wave (65) on different spatial variables $x_{1}$, $x_{2}$ and $x_{3}$ with respect to the time variable $t$ under some given parameters are shown in Figure 3 to demonstrate its propagation. Figure 3(a) illustrates the interaction of two bright breather waves on the periodic wave background, which is similar to that in Figure 1(a), while the two breather waves are characterized by different inverse widths $\bar{\mu}_{j}(j=1,\,2)$, propagation speeds $\bar{c}_{j}(j=1,\,2)$ and phase shifts $2\gamma_{j}(j=1,\,2)$, respectively. Similarly, in Figure 3(b), the two bright breather waves on periodic wave background propagate along the $t$ direction with the same inverse width $\omega_{2}$, and different phase shifts $2\gamma_{j}(j=1,\,2)$, and velocities are $\bar{c}_{j}/\bar{p}_{2,j}(j=1,\,2)$, respectively. The scenario depicted in Figure 3(c) bears a strong resemblance to that in Figure 3(b), and these parameters have different physical meanings in different contexts and are crucial for revealing the underlying physical mechanisms.
Figure 3: Propagation profiles and corresponding contour plots of the nonlinear wave solution (65) generated via the two-times Darboux transformation. The parameters are chosen as $k=0.3$, $\alpha=\pi/3$, $\beta=1$, $\lambda_{1}=1.8$, $\lambda_{2}=1.4$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, and $D=5$. (a) Evolution in the $x_{1}$$t$ plane with $x_{2}=x_{3}=1$; (b) evolution in the $x_{2}$$t$ plane with $x_{1}=x_{3}=1$; (c) evolution in the $x_{3}$$t$ plane with $x_{1}=x_{2}=1$.
Figure 4: Propagation profiles and corresponding contour plots of the nonlinear wave solution (67) generated via the two-times Darboux transformation. The parameters are chosen as $k=0.6$, $\alpha=\pi/3$, $\beta=1$, $\lambda_{1}=-2.1$, $\lambda_{2}=-2.3$, $\delta=1$, $\omega_{2}=\omega_{3}=1$, and $D=5$. (a) Evolution in the $x_{1}$$t$ plane with $x_{2}=x_{3}=1$; (b) evolution in the $x_{2}$$t$ plane with $x_{1}=x_{3}=1$; (c) evolution in the $x_{3}$$t$ plane with $x_{1}=x_{2}=1$.
@@
$\quad\ \ \times\Theta(\eta+\alpha_{1})\Theta(\eta+\alpha_{2})\mathrm{exp}(-\hat{\kappa}_{1}-\hat{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta+\alpha_{1})\Theta(\eta-\alpha_{2})\mathrm{exp}(-\hat{\kappa}_{1}+\hat{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta-\alpha_{1})\Theta(\eta+\alpha_{2})\mathrm{exp}(\hat{\kappa}_{1}-\hat{\kappa}_{2})$
@@
$\quad\ \ \times\Theta(\eta-\alpha_{1})\Theta(\eta-\alpha_{2})\mathrm{exp}(\hat{\kappa}_{1}+\hat{\kappa}_{2}),$
@@
$\quad\ \ \ \left.\times\sqrt{\lambda_{j}+1+\delta\omega_{2}}+D\right)/\hat{\mu}_{j},$
@@
[…]
@@
$\quad\ \ \times\Theta(\eta+\alpha_{2})\mathrm{exp}(-\bar{\kappa}_{1}-\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta-\alpha_{2})\mathrm{exp}(-\bar{\kappa}_{1}+\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta+\alpha_{2})\mathrm{exp}(\bar{\kappa}_{1}-\bar{\kappa}_{2})$
[…]
@@
$\quad\ \ \times\Theta(\eta-\alpha_{2})\mathrm{exp}(\bar{\kappa}_{1}+\bar{\kappa}_{2}),$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Figure 4 displays the evolution and corresponding contour plots for the second dark breather solution (67) across various spatial variables $x_{1}$, $x_{2}$, and $x_{3}$, in relation to the time variable $t$, for some specific parameter values. Figure 4 (a) depicts the interaction between two dark breather waves on the periodic background, distinguished by different inverse widths $\hat{\mu}_{j}(j=1,\,2)$, propagation speeds $\hat{c}_{j}(j=1,\,2)$, and phase shifts $2\alpha_{j}(j=1,\,2)$. In the same way, the interaction of two dark breather waves with the same inverse spatial width $\omega_{2}$ and different phase shifts $2\alpha_{j}(j=1,\,2)$ and velocities being $\hat{c}_{j}/\hat{p}_{2,j}~(j=1,\,2)$ on the periodic wave background is shown in Figure 4(b). The situation illustrated in Figure 4 (c) closely mirrors that in Figure 4 (b).